in one pot the ratio of milk and water is 3:5 and in another pot it is 6:1 in what ratio should the contents of the two pots be mixed so as to make the ratio between milk and water is to be 7:3

Answer by htmentor(1343)

Pot 1: Ratio M/W = 3/5 -> 3 parts milk, 5 parts water
Pot 2: Ratio M/W = 6/1 -> 6 parts milk, 1 part water
Let f = the amount of pot 1 used
Let g = the amount of pot 2 used
The mixture contains a ratio of milk to water of 7/3 -> 7 parts milk, 3 parts water
So we need 3f parts milk + 6g parts milk to make 7 parts milk
So we need 5f parts water + g parts water to make 3 parts water
3f + 6g = 7
5f + g = 3 -> g = 3 - 5f
3f + 6(3 - 5f) = 7
27f = 11 -> f = 11/27
Thus g = 3 - 5*11/27 = (81 - 55)/27 = 26/27
So f/g = (11/27)/(26/27) = 11/26
So pot 1 to pot 2 = 11:26